1. A Cylindrical capacitor having a length of 8 cm is made of two concentric rings with an inner radius as 3 cm and outer radius as 6 cm. Find the capacitance of the capacitor.




2. A parallel-plate capacitor is filled with polystyrene of 0.0002 m thickness. Find the plate area if the new capacitance after the insertion is 3.4 µF.

1
Expert's answer
2022-04-17T17:08:17-0400

1. The capacitance of a cylindrical capacitor is given as follows:


"C = \dfrac{2\pi \epsilon_0 L}{\ln(r_2/r_1)}"

where "L=8cm = 0.08m", "r_1 = 0.03m, r_2 = 0.06m" and "\epsilon_0 \approx 8.85\times 10^{-12} F/m" is vacuum permittivity. Thus, obtain:


"C = \dfrac{2\pi \cdot 8.85\times 10^{-12}\cdot 0.08}{\ln(0.06/0.03)} \approx 6.4\times 10^{-12}F"



2. The capacitance of a parallel plate capacitor is:


"C=\dfrac{\epsilon_0 \epsilon A}{d}"

where "A" is the plate area, "\epsilon =2.55" is the dielectric constant of polystyrene, and "d = 0.0002 m" is the distance between plates. Expressing "A", get:


"A = \dfrac{dC}{\epsilon_0 \epsilon} =\dfrac{0.0002m\cdot 3.4\times10^{-6}F }{8.85\times 10^{-12}F/m\cdot 2.55} \approx 30m^2"